OpenStudy (anonymous):

Jane poured 12 cups of water into a rectangular container measuring 12 inches by 8 inches by 9 centimeters what is the maximum volume of water that the rectangular container can hold? how many cubic inches of water was poured into the container? what was the height of the water in the container?

OpenStudy (igreen):

It wants you to find the volume of the container..but first we need to convert the measurements into similar measurements. 12 in, 8 in, 9 cm We can change 9 centimeters to approximately 3.54 inches. So we have: 12 in, 8 in, and 3.54 in. \(V = lwh\) \(V = (12)(8)(3.54)\) Can you multiply those three numbers together? @BugBear2015

OpenStudy (anonymous):

yeah im getting 339.84

OpenStudy (igreen):

Yep, that's the volume of the container in inches.

OpenStudy (igreen):

Err cubic inches.

OpenStudy (igreen):

How many cubic inches of water was poured into the container? Jane poured 12 cups of water in the container. Okay, 1 cup is about 14.44 cubic inches. So 12 cups will be \(12 \times 14.44\), can you multiply that?

OpenStudy (anonymous):

173.28?

OpenStudy (igreen):

Yep, so she poured in 173.28 cubic inches of water.

OpenStudy (igreen):

What was the height of the water in the container? To figure this out, we can divide the length times the width by the Volume of the water that was poured in. \(\dfrac{Volume}{Length \times Width}\) We already found the volume of the water, which was 173.28 cubic inches. And we are given the length and the width, 12 and 8 inches. \(\dfrac{173.28}{12 \times 8}\) Multiply: \(\dfrac{173.28}{96}\) Now can you divide that?

OpenStudy (anonymous):

1.805

OpenStudy (igreen):

Yep, so the height of the water in the container is 1.805 inches.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @BugBear2015

OpenStudy (anonymous):

thank you so much!